Find the equation of the normal to the solid 2x^2 −y^2 +8z^2 = 11 at a point where it intersects the line x−3 = z =(y+1)/2.
1
Expert's answer
2020-05-21T15:13:20-0400

The equation of the normal:

"\frac{x-x_0}{F'_x(M_0)}=\frac{y-y_0}{F'_y(M_0)}=\frac{z-z_0}{F'_z(M_0)}"

"F'_x=4x"

"F'_y=-2y"

"F'_z=16z"

Then:

"\frac{}{}" "\frac {x-x_0}{4x_0}=\frac {y-y_0}{-2y_0}=\frac {z-z_0}{16z_0}"

"4x_0=x_1-x_0"

"-2y_0=y_1-y_0"

"16z_0=z_1-z_0"

Where "(x_1,y_1,z_1)" is the point of intersection of the normal with the given line.

We have:

"x_1=5x_0, y_1=-y_0, z_1=15z_0"

So:

"5x_0-3=15z_0=\frac {-y_0+1}{2}"

"2x_0^2-y_0^2+8z_0^2=11"

"y_0=1-2(5x_0-3), z_0=\frac {5x_0-3}{15}"

"2x_0^2-(1-2(5x_0-3))^2+8(\frac {5x_0-3}{15})^2=11"

"2x_0^2-(1-4(5x_0-3)+4(25x_0^2-30x_0+9))+\frac {8}{225}(25x_0^2-30x_0+9)=11"

We have two cases:

"x_0=1.34, y_0=-6.4, z_0=0.25"

or

"x_0=0.08, y_0=6.2, z_0=-0.17"

Answer:

"\frac {x-1.34}{5.36}=\frac {y+6.4}{12.8}=\frac {z-0.25}{4}"

or

"\frac {x-0.08}{0.32}=\frac {y-6.2}{-12.4}=\frac {z+0.17}{-2.72}"